[figure: 9.7 cm]

Supertransferred Hyperfine Interactions in LaSrGa_{0.995}Cu_{0.005}O_{4}

Single Crystal and Isostructural Cuprates

Yu.V. Yablokov^{a}, O.A. Anikeenok^{b}, M.A. Augustyniak-Jabłokow^{a} and T.A. Ivanova^{c}.

aInstitute of Molecular Physics PAS, Smoluchowskiego 17, 60-179 Pozna, Poland

bKazan State University, Kremljevskaja 18, 420 008 Kazan, Russia

cKazan Physical Technical Institute RAN, Sibirskii Trakt 10/7, 420 029 Kazan, Russia

A nearly isotropic supertransferred hyperfine coupling of Cu^{2+} electron spin with the four neighboring Ga^{3+} nuclei in LaSrGa_{0.995}Cu_{0.005}O_{4}single crystal are discovered. EPR spectra parameters A_{Cu∥} = 143±.5; A_{Cu⊥} = 3±0.5; A_{Ga∥} = 18.4±0.3; A_{Ga⊥} 22.1±0.3 (in 10^{−4}cm^{−1}); g_{∥} = 2.320 ±0.005; g_{⊥} = 2.058 ±.005 are temperature independent in the 4.2 – 300 K range. An analysis of the obtained data is fulfilled in a framework of the configurational interaction theory. The cascade transfers of gallium 3s-and 3p-electron onto the copper orbitals through the 2s− and 2p−orbitals of the intermediate oxygens, due to the overlapping and covalence effects, are shown to explain the experimental parameters. The transfer integrals 3s, 3p (Ga) → unfilled orbital of oxygen are equal λ_{s} = 0.15, λ_{σ} = 0.18 and consistent with literature data for the iron group ions. Assuming identity of the on-set and supertransferred hyperfine fields in LaSrGa_{0.995}Cu_{0.005}O_{4}and in La_{2}CuO_{4} and using the results of NMR in the antiferromagnet La_{2}CuO_{4} [T.Tsuda et al., J.Phys.Soc.Jpn. 57, 2908(1989)] we estimate a zero-point spin deviation of ≈0 % in this compound. The shown comparability obtained local fields in LaSrGa_{0.995}Cu_{0.005}O_{4} with the corresponding fields in La_{2}CuO_{4} determined in the study of anisotropies of the Knight shifts and relaxation rates of the Cu nuclei and calculated by P. Hüsser et al. [Phys.Rev.61, 1567(2000)] is discussed.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Yu.V. Yablokov

Address for correspondence : Institute of Molecular Physics, PAS// ul. Smoluchowskiego 17//60-179 Poznań, POLAND

Email address : yablokov@ifmpan.poznan.pl